# The Prague Relational Learning Repository

### v2

@@ Abstract
The aim of the Prague Relational Learning Repository is to support machine learning research with multi-relational data. The repository currently contains 148 SQL databases hosted on a public MySQL server located at https://relational-data.org. The server is provided by getML to support the relational machine learning community (www.getml.com). A searchable meta-database provides metadata (e.g., the number of tables in the database, the number of rows and columns in the tables, the number of self-relationships).
@@ 4 Access and Contributions
relational
To contribute a database, please contact the repository administrators; a web contact form is available https://relational-data.org/contact. One possibility is to provide us with a MySQL dump of your database. Another option is to provide us with read access to your database on your server, so we can migrate the database to the public server. A web form for contributing is available  https://relational-data.org/contribute.
@@ 6 Conclusions
In this paper, we presented the Prague Relational Learning Repository (PRLR), an easily accessible collection of datasets for relational learning. The PRLR was designed with supervised learning in mind. To this end, the PRLR contains 148 ready to download datasets. One of the important features of the PRLR is that it provides meta-data about the datasets. The PRLR meta-data can be accessed at https://relational-data.org/.
@@ Acknowledgment
We are grateful to getML for sponsoring and maintaining the database server (www.getml.com). We would like to thank all of the donors who contributed data to the repository, and the author of the web page, Václav Ostrožlík. This work was supported by the Grant Agency of the Czech Technical University in Prague, grant No. SGS15/117/OHK3/1T/18.

### v3

@@ Abstract
The aim of the Prague Relational Learning Repository is to support machine learning research with multi-relational data. The repository currently contains 148 SQL databases hosted on a public MySQL server located at relational.fel.cvut.cz. The server is provided by CTU, the Czech technical university, to support the relational machine learning community. A searchable meta-database provides metadata (e.g., the number of tables in the database, the number of rows and columns in the tables, the number of self-relationships).
@@ 4 Access and Contributions
ctu-relational
To contribute a database, please contact the repository administrators; a web contact form is available https://relational.fel.cvut.cz/contact. One possibility is to provide us with a MySQL dump of your database. Another option is to provide us with read access to your database on your server, so we can migrate the database to the public server.
@@ 6 Conclusions
In this paper, we presented the Prague Relational Learning Repository (PRLR), an easily accessible collection of datasets for relational learning. The PRLR was designed with supervised learning in mind. To this end, the PRLR contains 148 ready to download datasets. One of the important features of the PRLR is that it provides meta-data about the datasets. The PRLR meta-data can be accessed at https://relational.fel.cvut.cz.
@@ Acknowledgment
We are grateful to CTU Prague for sponsoring and maintaining the database server. Our industry partner getML ensured service continuity by hosting the database server during a transition phase (www.getml.com). We would like to thank all of the donors who contributed data to the repository, and the author of the web page, Václav Ostrožlík. This work was supported by the Grant Agency of the Czech Technical University in Prague, grant No. SGS15/117/OHK3/1T/18.